4. Let X be continuous random variable whose PDF is given by for r>1 otherwise, where 0 is an unknown scalar parmeter. (a) Find the maximum likelihood estimator of . (b) Find a method-of-moments estimator of θ for the case when θ > 1. (c) Why can we not find a method-of-moments estimator when θ < 1? 151 151 151
Answer #1

(a)

The likelihood function for independent samples X1, X2, ..., Xn is,

-(0+1 for xi > 1

The log-likelihood function is,

mL(0) = n ln(0) _ (D+ 1) y ln(zi)

For maximum likelihood estimate,

rac{partial }{partial heta}lnL( heta) = n/ heta -sum_{i=1}^{n}~ln(x_i)= 0

Rightarrow heta = n/ sum_{i=1}^{n}~ln(x_i)

rac{partial^2 }{partial heta^2}lnL( heta) = -n/ heta^2 < 0

Thus, the maximum likelihood  estimate of heta is,

heta_{MLE} = n/ sum_{i=1}^{n}~ln(x_i)

(b)

First moment (mean) of f(x) is,

E(x) = int_{1}^{infty }xf(x)~dx = int_{1}^{infty } heta x^{- heta}~dx = heta left [x^{- heta + 1}/(- heta + 1) ight ]^{infty }_1 ----(1)

heta > 1 =>  - heta + 1 < 0 => x^{- heta + 1} ightarrow 0 as x ightarrow infty .

E(x) = heta left [x^{- heta + 1}/(- heta + 1) ight ]^{infty }_1 = heta / ( heta -1)

Thus, sum_{i=1}^{n} X_i /n= ar{X} = E(x) = heta / ( heta -1)

Rightarrow ar{X} ( heta -1) = heta

Rightarrow ar{X} heta -ar{X} = heta

Rightarrow ar{X} heta - heta = ar{X}

Rightarrow heta(ar{X} -1) = ar{X}

Rightarrow heta = ar{X} /(ar{X} -1)

So, the method of moments estimator is,

heta_{MOM} = ar{X} /(ar{X} -1) where ar{X} = sum_{i=1}^{n} X_i /n

(c)

For heta = 1, the first moment (mean) of f(x,) E(x) = heta / ( heta -1) is undefined.

For heta < 1,

- heta + 1 > 0 => x^{- heta + 1} ightarrow infty as x ightarrow infty

So, by equation (1), the first moment (mean) of f(x) is undefined for heta < 1. Hence we cannot find a method of moments estimator for heta le 1.